# Improved Scaling Laws in Linear Regression via Data Reuse

### v1

@@ 7 Conclusion
Many directions remain open for future research. First, our analysis focuses on multi-pass SGD with batch size one; it would be interesting to understand how the test error scales with the batch size and to develop corresponding batch size scaling laws (see Jain et al., 2017). Another important direction is to study how data reuse interacts with other optimization algorithms, such as SGD with momentum and Adam. In addition, it would be valuable to extend our analysis to non-linear settings and classification problems, such as logistic regression, kernel methods, and neural networks. Notably, modern large language model pretraining is based on minimizing the cross-entropy loss for next-word prediction. Understanding the scaling behavior in logistic regression—the simplest classification model—thus represents an important step toward unraveling the mysteries of LLM scaling.
@@ A.1 Comments and additional notation
Throughout the appendix, we assume without loss of generality that the covariance matrix $\mathbf{H}$ is diagonal, with diagonal entries given by the eigenvalues $(\lambda_{i})_{i\geq 1}$ in non-increasing order. This reduction is justified by the rotational invariance of the Gaussian sketching matrix $\mathbf{S}$. Under this diagonalization, Assumption 1D can be restated more explicitly as follows:
@@
for $\text{for all }i\neq j,\quad{\mathbb{E}}[\mathbf{w}^{*}_{i}\mathbf{w}^{*}_{j}]=0;\qquad\text{and for all }i>0,\quad{\mathbb{E}}[\lambda_{i}\mathbf{w}_{i}^{*2}]\eqsim i^{-b},\quad\text{for some }b>1.$
@@ A.1.1 Assumptions on the stepsize
Under the notations in Theorem 3.1 and its proof, with probability at least $1-\exp(-\Omega(M))$ over the randomness of $\mathbf{S}$, we have
@@
=$\overset{(i)}{=}\sigma^{2}+\mathbb{E}\big{[}\langle\mathbf{x},\mathbf{S}^{\top}\mathbf{v}^{*}-\mathbf{w}^{*}\rangle^{2}\big{]}+\mathbb{E}\big{[}\langle\mathbf{x},\mathbf{S}^{\top}(\mathbf{v}_{L}-\mathbf{v}^{*})\rangle^{2}\big{]}$
=$\overset{(ii)}{=}{\sigma^{2}}+{\mathbb{E}\big{[}\langle\mathbf{x},\mathbf{S}^{\top}\mathbf{v}^{*}-\mathbf{w}^{*}\rangle^{2}\big{]}}+{\mathbb{E}\big{[}\langle\mathbf{x},\mathbf{S}^{\top}(\bm{\theta}_{L}-\mathbf{v}^{*})\rangle^{2}\big{]}}+{\mathbb{E}\big{[}\langle\mathbf{x},\mathbf{S}^{\top}(\mathbf{v}_{L}-\bm{\theta}_{L})\rangle^{2}\big{]}}$
@@
We have from the proof of Theorem 4.2 in Lin et al. (2024) that ${\mathbb{E}}_{\mathbf{w}^{*}}\mathsf{Approx}={\mathbb{E}}_{\mathbf{w}^{*}}\|\mathbf{S}^{\top}\mathbf{v}^{*}-\mathbf{w}^{*}\|_{\mathbf{H}}^{2}\eqsim M^{1-b}$ with probability at least $1-e^{-\Omega(M)}$ over $\mathbf{S}.$
@@
Let $\tilde{\epsilon}_{i}=y_{i}-\mathbf{x}_{i}^{\top}\mathbf{S}^{\top}\mathbf{v}^{*}$ for $i\in[N]$ and write $\tilde{\bm{\epsilon}}=(\tilde{\epsilon}_{1},\ldots,\tilde{\epsilon}_{N})^{\top}$. It can be verified that, conditioned on $\mathbf{S}$, ${\mathbb{E}}[\tilde{\epsilon}_{i}]=0$ and $\tilde{\epsilon}_{i}$ is independent of $\mathbf{S}\mathbf{x}_{i}$. Moreover,
@@
$=\sigma^{2}+{\mathbb{E}}_{\mathbf{w}^{*}}[\mathbf{w}^{*\top}\mathbf{H}^{1/2}(\mathbf{I}-\mathbf{H}^{1/2}\mathbf{S}^{\top}(\mathbf{S}\mathbf{H}\mathbf{S}^{\top})^{-1}\mathbf{H}^{1/2})\mathbf{H}^{1/2}\mathbf{w}^{*}]$
@@
$\bm{\theta}_{t}-\mathbf{v}^{*}$ $=\bm{\theta}_{t}-\mathbf{v}^{*}-\frac{\gamma_{t}}{N}\mathbf{S}\mathbf{X}^{\top}(\mathbf{X}\mathbf{S}^{\top}\bm{\theta}_{t-1}-\mathbf{y})=\bm{\theta}_{t-1}-\mathbf{v}^{*}-\frac{\gamma_{t}}{N}\mathbf{S}\mathbf{X}^{\top}(\mathbf{X}\mathbf{S}^{\top}(\bm{\theta}_{t-1}-\mathbf{v}^{*})-\tilde{\bm{\epsilon}})$
@@ B.1 An upper bound
Define $\mathsf{M}:=\prod_{t=1}^{L}(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}}){\bm{\Sigma}}\prod_{t=1}^{L}(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}})$ and recall that ${\mathbf{v}^{*}}=\big{(}\mathbf{S}\mathbf{H}\mathbf{S}^{\top}\big{)}^{-1}\mathbf{S}\mathbf{H}{\mathbf{w}^{*}}$. Substituting
@@
$\prod_{t=1}^{L}(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}})$ $=\prod_{t=1}^{L-1}(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}})-\gamma_{L}{\bm{\widehat{\Sigma}}}\prod_{t=1}^{L-1}(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}})\overset{(iii)}{=}\prod_{t=1}^{L-1}(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}})-\frac{1}{N}\mathbf{S}\mathbf{X}^{\top}\Big{[}\gamma_{L}\prod_{t=1}^{L-1}(\mathbf{I}-\gamma_{t}\bm{\overline{\Sigma}})\Big{]}\mathbf{X}\mathbf{S}^{\top}$
@@
[…]
@@
$\mathbf{Z}{\bm{\Sigma}}^{2}\mathbf{Z}^{\top}\preceq 5\cdot(\mathbf{Z}{\bm{\Sigma}}\mathbf{Z}^{\top})^{2}$
@@ B.2 A lower bound
with probability at least $1-e^{-\Omega(M)}$, where $\overline{t}:={\mathbb{E}}_{\mathbf{X}}[\#\{i\in[M]:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}>1/4\}]$ and $(\widehat{\lambda}_{i})_{i=1}^{M}$ are the eigenvalues of ${\bm{\widehat{\Sigma}}}$.
@@
Let $\mathsf{C}_{L}:=\prod_{t=1}^{L}(\mathbf{I}-\gamma_{t}{\bm{\widehat{\Sigma}}})$. By definition,
@@
[…]
@@
Let ${\bm{\widehat{\Sigma}}}=\widehat{\mathbf{U}}\widehat{\bm{\Gamma}}\widehat{\mathbf{U}}^{\top}$ be the singular value decomposition with $\widehat{\bm{\Gamma}}=\operatorname{diag}\{\widehat{\lambda}_{1},\ldots,\widehat{\lambda}_{M}\}$ and $\widehat{\lambda}_{1}\geq\ldots\geq\widehat{\lambda}_{M}$. Then we have

### v2

@@ 7 Conclusion
Our analysis has several limitations. One limitation is the assumption that the eigenvectors of the prior and data covariance are aligned (implied by Assumption 1D). While this assumption cannot be fully removed without affecting the error rate, it would be interesting to investigate what alternative rates are achieved when the eigenvectors are not aligned. Another limitation is that our lower bound results require Gaussian design of the covariates (i.e., Assumption 1A); a next step is to extend them to non-Gaussian design.
Many other directions remain open for future research. First, our analysis focuses on multi-pass SGD with batch size one; it would be interesting to understand how the test error scales with the batch size and to develop corresponding batch size scaling laws (see Jain et al., 2017). Another important direction is to study how data reuse interacts with other optimization algorithms, such as SGD with momentum or $\ell_{2}$-regularization and Adam. In addition, it would be valuable to extend our analysis to non-linear settings and classification problems, such as logistic regression, kernel methods, and neural networks. Notably, modern large language model pretraining is based on minimizing the cross-entropy loss for next-word prediction. Understanding the scaling behavior in logistic regression—the simplest classification model—thus represents an important step toward unraveling the mysteries of LLM scaling.
@@ A.1 Comments and additional notation
Throughout the appendix (except for Appendix A.3), we assume without loss of generality that the covariance matrix $\mathbf{H}$ is diagonal, with diagonal entries given by the eigenvalues $(\lambda_{i})_{i\geq 1}$ in non-increasing order. This reduction is justified by the rotational invariance of the Gaussian sketching matrix $\mathbf{S}$. Under this diagonalization, Assumption 1D can be restated more explicitly as follows:
@@
$\text{for all }i\neq j,\quad{\mathbb{E}}[\mathbf{w}^{*}_{i}\mathbf{w}^{*}_{j}]=0;\qquad\text{and for all }i>0,\quad{\mathbb{E}}[\lambda_{i}\mathbf{w}_{i}^{*2}]\eqsim i^{-b},\quad\text{for some }b>1.$
@@ A.1.1 Assumptions on the stepsize
Under the notations in Theorem 3.1 and its proof, assume that, with probability at least $1-\exp(-\Omega(M))$ over the randomness of $\mathbf{S}$,
@@
$\overset{(i)}{=}\sigma^{2}+\mathbb{E}\big[\langle\mathbf{x},\mathbf{S}^{\top}\mathbf{v}^{*}-\mathbf{w}^{*}\rangle^{2}\big]+\mathbb{E}\big[\langle\mathbf{x},\mathbf{S}^{\top}(\mathbf{v}_{L}-\mathbf{v}^{*})\rangle^{2}\big]$
$\overset{(ii)}{=}{\sigma^{2}}+{\mathbb{E}\big[\langle\mathbf{x},\mathbf{S}^{\top}\mathbf{v}^{*}-\mathbf{w}^{*}\rangle^{2}\big]}+{\mathbb{E}\big[\langle\mathbf{x},\mathbf{S}^{\top}(\bm{\theta}_{L}-\mathbf{v}^{*})\rangle^{2}\big]}+{\mathbb{E}\big[\langle\mathbf{x},\mathbf{S}^{\top}(\mathbf{v}_{L}-\bm{\theta}_{L})\rangle^{2}\big]}$
@@
We have from Lemma C.5 in Lin et al. (2024) that ${\mathbb{E}}_{\mathbf{w}^{*}}\mathsf{Approx}={\mathbb{E}}_{\mathbf{w}^{*}}\|\mathbf{S}^{\top}\mathbf{v}^{*}-\mathbf{w}^{*}\|_{\mathbf{H}}^{2}\eqsim M^{1-b}$ with probability at least $1-e^{-\Omega(M)}$ over $\mathbf{S}.$
@@
Let $\tilde{\epsilon}_{i}=y_{i}-\mathbf{x}_{i}^{\top}\mathbf{S}^{\top}\mathbf{v}^{*}$ for $i\in[N]$ and write $\tilde{\bm{\epsilon}}=(\tilde{\epsilon}_{1},\ldots,\tilde{\epsilon}_{N})^{\top}$. It can be verified that, conditioned on $(\mathbf{S},\mathbf{w}^{*})$, ${\mathbb{E}}[\tilde{\epsilon}_{i}]=0$ and $\tilde{\epsilon}_{i}$ is independent of $\mathbf{S}\mathbf{x}_{i}$. Moreover,
@@
$=\sigma^{2}+{\mathbb{E}}_{\mathbf{w}^{*}}[\mathbf{w}^{*\top}\mathbf{H}^{1/2}(\mathbf{I}-\mathbf{H}^{1/2}\mathbf{S}^{\top}(\mathbf{S}\mathbf{H}\mathbf{S}^{\top})^{-1}\mathbf{S}\mathbf{H}^{1/2})\mathbf{H}^{1/2}\mathbf{w}^{*}]$
@@
[…]
@@
A.3 Relaxation of Assumption 1
In this section, we show that some conditions in Assumption 1 can be further relaxed. Concretely, we have
  • (a).
    The exact alignment of the eigenvectors of the prior and data covariance in Assumption 1D is not necessary. All results in Section 3 remain valid if Assumption 1D is replaced by
    Assumption 1D’ (Approximate source condition).
    Let $(\lambda_{i},\mathbf{v}_{i})_{i>0}$ be the eigenvalues and eigenvectors of $\mathbf{H}$ and let $\mathbf{H}^{\mathbf{w}}={\mathbb{E}}[\mathbf{w}^{*}\mathbf{w}^{*\top}]$. Assume $c\widetilde{\mathbf{H}}^{\mathbf{w}}\preceq\mathbf{H}^{\mathbf{w}}\preceq c^{\prime}\widetilde{\mathbf{H}}^{\mathbf{w}}$ for some absolute constants $c^{\prime}\geq c>0$ and $\widetilde{\mathbf{H}}^{\mathbf{w}}\succeq 0$ such that
[…]
  • (b).
    To establish the upper bounds for $\mathsf{Bias},\mathsf{Var},\mathsf{Approx}$ in Theorem 3.1 and the upper bounds in Corollary 3.23.4, Assumption 1A can be relaxed to
    Assumption 1A’ (sub-Gaussian design).
    $\mathbf{x}=\mathbf{H}^{1/2}\widetilde{\mathbf{x}}$, where ${\mathbb{E}}[{\widetilde{\mathbf{x}}}{\widetilde{\mathbf{x}}}^{\top}]=\mathbf{I}$, and the vector $\widetilde{\mathbf{x}}$} is symmetric and 1-sub-Gaussian, i.e., ${\widetilde{\mathbf{x}}}\overset{d}{=}-{\widetilde{\mathbf{x}}}$ and ${\mathbb{E}}[e^{\lambda\langle\mathbf{v},{\widetilde{\mathbf{x}}}\rangle}]\leq e^{\lambda^{2}/2}$ for any unit vector $\mathbf{v}$ and all $\lambda\in{\mathbb{R}}.$
We provide some justification of the two relaxations below.
Justification of (a).
By checking the proof of Theorem 3.1 and its corollaries, it can be seen that Assumption 1D is used to (1) give matching upper and lower bounds on ${\mathbb{E}}_{\mathbf{w}^{*}}[\|\mathbf{w}_{{0:k}}^{*}\|^{2}_{2}],{\mathbb{E}}_{\mathbf{w}^{*}}[\|\mathbf{w}_{{k:\infty}}^{*}\|^{2}_{\mathbf{H}_{k:\infty}}],\mu_{i}(\mathbf{S}\mathbf{H}\mathbf{H}^{\mathbf{w}}\mathbf{H}\mathbf{S}^{\top})$ for any $k\geq 0$ and $i\in[M]$ when controlling the approximation and bias error (see Lemma C.5 in Lin et al. (2024) and Lemma B.3); (2) give matching upper and lower bounds on ${\mathbb{E}}[\|\mathbf{w}^{*}\|^{2}_{\mathbf{H}}]$ when controlling the fluctuation error (see Lemma D.5). Under the alternative Assumption 1D’, it is readily verified that the same bounds on these quantities can be established up to constant factors. Concretely, suppose there exists some parameter $\widetilde{\mathbf{w}}^{*}$ with prior ${\mathbb{E}}[\widetilde{\mathbf{w}}^{*}\widetilde{\mathbf{w}}^{*\top}]=\widetilde{\mathbf{H}}^{\mathbf{w}}$. Then $\widetilde{\mathbf{w}}^{*}$ satisfies Assumption 1D and
${\mathbb{E}}_{\mathbf{w}^{*}}[\|\mathbf{w}_{{0:k}}^{*}\|^{2}_{2}]=\operatorname{tr}(\mathbf{H}^{\mathbf{w}}_{0:k})$ $\eqsim\operatorname{tr}(\widetilde{\mathbf{H}}^{\mathbf{w}}_{{0:k}})={\mathbb{E}}_{\widetilde{\mathbf{w}}^{*}}[\|\widetilde{\mathbf{w}}_{{0:k}}^{*}\|^{2}_{2}],$
${\mathbb{E}}_{\mathbf{w}^{*}}[\|\mathbf{w}_{{k:\infty}}^{*}\|^{2}_{\mathbf{H}_{k:\infty}}]=\operatorname{tr}(\mathbf{H}_{k:\infty}^{1/2}\mathbf{H}^{\mathbf{w}}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2})$ $\eqsim\operatorname{tr}(\mathbf{H}_{k:\infty}^{1/2}\widetilde{\mathbf{H}}^{\mathbf{w}}_{{k:\infty}}\mathbf{H}_{k:\infty}^{1/2})={\mathbb{E}}_{\widetilde{\mathbf{w}}^{*}}[\|\widetilde{\mathbf{w}}_{{k:\infty}}^{*}\|^{2}_{\mathbf{H}_{k:\infty}}],$
$\mu_{i}(\mathbf{S}\mathbf{H}\mathbf{H}^{\mathbf{w}}\mathbf{H}\mathbf{S}^{\top})$ $\overset{(i)}{\eqsim}\mu_{i}(\mathbf{S}\mathbf{H}\widetilde{\mathbf{H}}^{\mathbf{w}}\mathbf{H}\mathbf{S}^{\top}),$
${\mathbb{E}}[\|\mathbf{w}^{*}\|^{2}_{\mathbf{H}}]=\operatorname{tr}(\mathbf{H}^{1/2}\mathbf{H}^{\mathbf{w}}\mathbf{H}^{1/2})$ $\eqsim\operatorname{tr}(\mathbf{H}^{1/2}\widetilde{\mathbf{H}}^{\mathbf{w}}\mathbf{H}^{1/2})={\mathbb{E}}_{\widetilde{\mathbf{w}}^{*}}[\|\widetilde{\mathbf{w}}^{*}\|^{2}_{\mathbf{H}}],$
where step (i) follows from the fact that $\mu_{i}(\mathbf{A})\leq\mu_{i}(\mathbf{B})$ for all $i$ and any $\bm{0}\preceq\mathbf{A}\preceq\mathbf{B}$. Therefore, the proof of Theorem 3.1 and its corollaries goes through under the alternative Assumption 1D’.
Justification of (b).
In short, the relaxation can be made since the Gaussian assumption is mainly used to establish certain concentration bounds (e.g., Bernstein’s inequality), which also hold for sub-Gaussian vectors. More specifically, the Gaussian design in Assumption 1A’ is used in our proof mainly in three ways: (1) to establish concentration bounds on the sample covariance (e.g., Eq. 10); (2) to allow the use of technical lemmas in Appendix E (e.g., Lemma E.3 and E.4); (3) to control the norm of sketched samples (e.g., to control $B_{\bm{\nu}}$ in Eq. 20).
Correspondingly, when $\mathbf{x}$ satisfies the alternative Assumption 1A’, we can show that (1) the same concentration bounds hold on the sub-Gaussian sample covariance by e.g., Theorem 6.5 in Wainwright (2019); (2) all technical lemmas in Appendix E hold when the Gaussian sketching $\mathbf{S}$ is replaced by a row-wise sub-Gaussian matrix by concentration bounds on quadratic forms of sub-Gaussian vectors (e.g., Theorem 1 in Hsu et al. (2012)), and on sub-Gaussian covariance matrices (e.g., Example 1.5 in Zhivotovskiy (2024)); (3) the norm of sketched samples satisfy the same concentration bounds by e.g., Theorem 6.5 in Wainwright (2019). Moreover, the symmetry assumption $\widetilde{\mathbf{x}}\overset{d}{=}-\widetilde{\mathbf{x}}$ is made to ensure that the proof of the commutativity result in Eq. (15a) remains valid.
On the other hand, for the lower bounds, the Gaussian assumption is still required in order to establish the conditional independence of $\tilde{\epsilon}_{i}=y_{i}-\mathbf{x}_{i}^{\top}\mathbf{S}^{\top}\mathbf{v}^{*}$ and $\mathbf{S}\mathbf{x}_{i}$ given $(\mathbf{S},\mathbf{w}^{*})$ in the proof of Theorem 3.1 and Lemma D.4.
In addition, we also conduct experiments to check our jusification of (b). We generate data $\mathbf{x}=(x_{1},\ldots,x_{d})^{\top}$ from the distribution where $x_{i}$ are independent and
$\mathbb{P}(x_{i}=1)=\mathbb{P}(x_{i}=-1)=i^{-a}/c_{0},~~~\mathbb{P}(x_{i}=0)=1-2\cdot\mathbb{P}(x_{i}=1),$
with $a=2,b=1.5$ and $c_{0}={2\sum_{i=1}^{d}i^{-a}}$. Note that $\mathbf{x}$ satisfies Assumption 1A’ but not Assumption 1A when $d=\infty$. We run the experiment under the same setting and choice of hyperparameters as in Figure 1(a). Similar to the Gaussian case, in Figure 2, we observe that the excess test error of one-pass SGD and multi-pass SGD both exhibit power-law scaling in the number of effective samples $N_{\mathtt{eff}}$}. Moreover, the fitted slopes are both close to the theoretical prediction in Corollary 3.3 ($0.34\approx 0.33=(1-b)/b$ for multi-pass SGD and $0.26\approx 0.25=(1-b)/a$ for one-pass SGD).
Figure 2: Multi-pass SGD versus one-pass SGD for non-Gaussian design. Multi-pass SGD is ran for $L\eqsim N^{a/b}$ steps. We use linear functions to fit the excess test error in log-log scale. The fitted exponents ($k$) are close to the theoretical predictions in Corollary 3.3. The errorbars denote the $\pm 1$ standard deviation of the expected excess test error over 100100 i.i.d. samples of $(\mathbf{S},\mathbf{w}^{*})$. Parameters: $\sigma^{2}=1$, $d=10000$, $M=1000$, $\gamma=0.1.$
@@ B.1 An upper bound
Without loss of generality, we assume the covariance matrix $\mathbf{H}=\operatorname{diag}\{\lambda_{1},\lambda_{2},\ldots,\lambda_{d}\}$ where $\lambda_{i}\geq\lambda_{j}$ for any $i\geq j$. Let $(\tilde{\lambda}_{1},\tilde{\lambda}_{2},\ldots,\tilde{\lambda}_{M})$ denote the eigenvalues of $\bm{\Sigma}$} in non-increasing order. Moreover, we introduce $\mathbf{z}_{1},\ldots\mathbf{z}_{N}\overset{iid}{\sim}\mathcal{N}(0,\mathbf{I}_{M}/N)$ and write $\mathbf{Z}=(\mathbf{z}_{1},\ldots,\mathbf{z}_{N})^{\top}$. It can be verified that $\mathbf{X}\mathbf{S}^{\top}/\sqrt{N}\overset{d}{=}\mathbf{Z}{\bm{\Sigma}}^{1/2}$ conditioned on $\mathbf{S}$. Throughout the proof, by a union bound argument, we w.l.o.g. assume the conditions (1), (2), (3) and (4) in Assumption 3 always hold.
@@
[…]
@@
where step (iii) uses $\mathbf{X}\mathbf{S}^{\top}(\mathbf{I}-\gamma_{L}{\bm{\widehat{\Sigma}}})=(\mathbf{I}-\gamma_{L}\bm{\overline{\Sigma}})\mathbf{X}\mathbf{S}^{\top}.$ Recall that $\mathbf{X}\mathbf{S}^{\top}/\sqrt{N}\overset{d}{=}\mathbf{Z}{\bm{\Sigma}}^{1/2}$ conditioned on $\mathbf{S}$, we can thus rewrite
@@
$\mathbf{Z}{\bm{\Sigma}}^{2}\mathbf{Z}^{\top}\preceq 5\cdot(\mathbf{Z}{\bm{\Sigma}}\mathbf{Z}^{\top})^{2}$ (10)
@@ B.2 A lower bound
with probability at least $1-e^{-\Omega(M)}$, where $\overline{t}:={\mathbb{E}}_{\mathbf{X}}[\#\{i\in[M]:\widehat{\lambda}_{i}{L_{\mathtt{eff}}}\gamma_{0}>1/4\}]$ and $(\widehat{\lambda}_{i})_{i=1}^{M}$ are the eigenvalues of $^{\bm{\widehat{\Sigma}}}$.
@@
Similar to the proof of Lemma B.1, w.l.o.g., we assume the covariance matrix $\mathbf{H}=\operatorname{diag}\{\lambda_{1},\lambda_{2},\ldots\\ ,\lambda_{d}\}$ where $\lambda_{i}\geq\lambda_{j}$ for any $i\geq j$. Let $(\tilde{\lambda}_{1},\tilde{\lambda}_{2},\ldots,\tilde{\lambda}_{M})$ denote the eigenvalues of $\bm{\Sigma}$} in non-increasing order. Moreover, we introduce $\mathbf{z}_{1},\ldots\mathbf{z}_{N}\overset{iid}{\sim}\mathcal{N}(0,\mathbf{I}_{M}/N)$ and write $\mathbf{Z}=(\mathbf{z}_{1},\ldots,\mathbf{z}_{N})^{\top}$. It can be shown that $\mathbf{X}\mathbf{S}^{\top}/\sqrt{N}\overset{d}{=}\mathbf{Z}{\bm{\Sigma}}^{1/2}$ conditioned on $\mathbf{S}$.
@@
Define the learning rate $\gamma_{0}(\Gamma)=\min\{1/[4\max_{j}\|\Gamma_{\cdot,j}\|_{2}^{2}],\gamma\}$ for any matrix $\Gamma\in{\mathbb{R}}^{M\times N}$ and define $\gamma_{t}(\Gamma)$ for all $t\in[L]$ according to (2). Let ${\bm{\Sigma}}=\mathbf{U}\tilde{\bm{\Gamma}}\mathbf{U}^{\top}$ be the singular value decomposition with $\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{M}$ and $\tilde{\bm{\Gamma}}=\operatorname{diag}\{\tilde{\lambda}_{1},\ldots,\tilde{\lambda}_{M}\}$ being a diagonal matrix with $\tilde{\lambda}_{1}\geq\tilde{\lambda}_{2}\geq\ldots\tilde{\lambda}_{M}\geq 0$. Note that $\mathbf{S}\mathbf{X}^{\top}/\sqrt{N}\overset{d}{=}\mathbf{U}\tilde{\bm{\Gamma}}^{1/2}\mathbf{Z}^{\top}$ conditioned on $\mathbf{S}$ and $\mathbf{U}^{\top}{\bm{\widehat{\Sigma}}}\mathbf{U}\overset{d}{=}\tilde{\bm{\Gamma}}^{1/2}\mathbf{Z}^{\top}\mathbf{Z}\tilde{\bm{\Gamma}}^{1/2}$. Therefore
@@
Let $^{\bm{\widehat{\Sigma}}}=\widehat{\mathbf{U}}\widehat{\bm{\Gamma}}\widehat{\mathbf{U}}^{\top}$ be the singular value decomposition with $\widehat{\bm{\Gamma}}=\operatorname{diag}\{\widehat{\lambda}_{1},\ldots,\widehat{\lambda}_{M}\}$ and $\widehat{\lambda}_{1}\geq\ldots\geq\widehat{\lambda}_{M}$. Then we have